OpenStudy (anonymous):

If one endpoint of AD is A(-3,-8) and the midpoint of AD is M(3,-4) what is the other endpoint D?

OpenStudy (anonymous):

to help you answer your question ill ask you a question. When you draw out the points how many points is it from A (-3,-8) to Point M(3,-4)

OpenStudy (anonymous):

you'll continue that many points in the same direction to get the Endpoint D

OpenStudy (anonymous):

Is there a equation i can use or do i just map it out?

OpenStudy (anonymous):

mapping it out shouldnt be tough

OpenStudy (anonymous):

just use a piece of graph paper (:

OpenStudy (anonymous):

you can use this if yu dont have graph paper http://www.garrettbartley.com/graphpaper.html

OpenStudy (anonymous):

i got you rise 4 and slide 6:)

OpenStudy (anonymous):

thats not a option on my choices tho its.. A 3.-4 B-7,11 C 7,-6 D 9,0

OpenStudy (anonymous):

this is so confusing xD HAHA

OpenStudy (anonymous):

you need to put the actual coordinates you got at the D Endpoint i think

OpenStudy (anonymous):

Huhhh?

OpenStudy (jdoe0001):

you know the midpoint formula?

OpenStudy (jdoe0001):

\(\bf \textit{middle point of 2 points }\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) A&({\color{red}{ -3}}\quad ,&{\color{blue}{ -8}})\quad % (c,d) D&({\color{red}{ \square }}\quad ,&{\color{blue}{ \square }}) \end{array}\qquad % coordinates of midpoint \left(\cfrac{{\color{red}{ x_2}} + {\color{red}{ x_1}}}{2}\quad ,\quad \cfrac{{\color{blue}{ y_2}} + {\color{blue}{ y_1}}}{2} \right)=M\) follow that? if you dunno the midpoint formula... you may want to refresh on that one first

OpenStudy (jdoe0001):

we now what A coordinates are we dunno what D coordinates we also know that the coordinates for M are so we can use that and say that \(\bf \textit{middle point of 2 points }\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) A&({\color{red}{ -3}}\quad ,&{\color{blue}{ -8}})\quad % (c,d) D&({\color{red}{ \square }}\quad ,&{\color{blue}{ \square }}) \end{array}\qquad % coordinates of midpoint \left(\cfrac{{\color{red}{ \square }} + {\color{red}{ (-3)}}}{2}\quad ,\quad \cfrac{{\color{blue}{ \square }} + {\color{blue}{ (-8)}}}{2} \right)=M \\ \quad \\ M=(3,-4)\qquad thus \\ \quad \\ \left(\cfrac{{\color{red}{ \square }} + {\color{red}{ (-3)}}}{2}\quad ,\quad \cfrac{{\color{blue}{ \square }} + {\color{blue}{ (-8)}}}{2} \right)=M \\ \quad \\ \left(\cfrac{{\color{red}{ \square }} + {\color{red}{ (-3)}}}{2}\quad ,\quad \cfrac{{\color{blue}{ \square }} + {\color{blue}{ (-8)}}}{2} \right)=(3,-4)\implies \begin{cases} \cfrac{{\color{red}{ \square }} + {\color{red}{ (-3)}}}{2}=3 \\ \quad \\ \cfrac{{\color{blue}{ \square }} + {\color{blue}{ (-8)}}}{2}=-4 \end{cases}\) so just solve for the ones you don't have

OpenStudy (jdoe0001):

and recall that D is \(\bf ({\color{red}{ \square }}\quad ,{\color{blue}{ \square }})\) so.. whatever those variables are, is D coordinates